# Can AI Dream of Unseen Galaxies? Conditional Diffusion Model for Galaxy Morphology Augmentation

### v1

@@ 4.3 Enhancing Early-type Dust Lane Galaxy Detection via Extrapolation on Rare Morphologies
Figure 6: Generative extrapolation of early-type galaxies with prominent dust lanes (D-ETGs). While early-type galaxies are commonly quiescent, our generative model synthesizes D-ETGs by merging the morphology feature in early-type galaxy and star-forming feature (dust lane) in late-type galaxies. Both early-type galaxies and dust lanes are common in the universe, but early-type galaxies with dust lanes are rare since they are origin from recent merging events. Our model learns these two concepts from extensive datasets and extrapolates them to synthesize rare D-ETGs. These synthesized D-ETG images are then used as training data for classifiers, facilitating the identification of 520 D-ETGs within a sample of 239,695 GZ2 images as shown in Figure 7.
@@
Figure 7: Examples of D-ETGs synthesized by our diffusion model (upper panel) and newly identified instances using machine learning models augmented with the synthesized data (lower panel). Our diffusion model generates high-fidelity D-ETG images by prompts like “sdss, {completely round:1.2} galaxy, {dust lane:1.3}” or “sdss, in-between round galaxy, {dust lane:1.3}”, and these synthesized images enhance the efficiency of identifying such objects in real datasets. We manually assign weights to different morphology features through cross-attention mechanism to emphasize generation of desired features as in Section 2.
@@
We initiate our investigation with an experimental test. Using the D-ETG catalog from Kaviraj et al. (2012), we aim to verify the effectiveness of our pipeline. In the experimental setup, we randomly select 100 galaxies from the D-ETG catalog in Kaviraj et al. (2012) and 10,000 galaxies from the GZ2 dataset to construct the training set, composing a experimental dataset with 1/100 postive/negative sample ratios, mimic the data distribution in Kaviraj et al. (2012). Similar to the approach in Section 4.2, we use pure supervised learning without any additional adjustments as our baseline.
Again, we introduce Focal Loss as an alternative basedline for dealing with imbalance classes as in Section 4.2. This method reweights the contributions of positive and negative samples in the loss function. We note that we defer detailed comparisons with more advanced few - shot algorithms and unsupervised classification algorithms to future studies, as their inclusion is beyond the scope of this work. More importantly, our methodologies are not in competition with these algorithms but can be used together to enhance these algorithms through data augmentation, a topic that we will discuss further in Section 5.
@@
Figure 8: Kernel density estimates (KDE) of redshift distributions for D-ETGs identified by our model (red) and Kaviraj et al. (2012) (blue). Redshift values are obtained from the GSWLC Salim et al. (2016). Our sample encompasses relatively higher redshifts, a regime that is more challenging for human volunteers to identify, thereby demonstrating the efficacy of our synthesized image-enhanced machine learning model in detecting rare galaxy populations.
@@
Here, we apply our method to 239,695 Galaxy Zoo 2 (GZ2) images to identify previously undetected D-type early-type galaxies (D-ETGs). We first construct training data: 10,000 synthetic D-ETGs as positive samples and 10,000 randomly selected normal galaxies as negative samples. Given the low occurrence of D-ETGs, we assume minimal contamination in the negative sample set. We follow the training procedure described in Section 4.2. After the initial training, the model was applied to the entire GZ2 dataset, identifying 8,184 positive candidates. These candidates were then visually inspected, and the classifier was retrained based on the validated results, yielding 2,488 refined positive samples. This iterative process was repeated once more, resulting in an additional 1,062 positive samples. All candidate samples in the three iteration training stages were visually validated by three authors (CM, ZS, and TJ). Final selection required agreement from at least two reviewers, resulting in a catalog of 520 high-confidence D-ETGs overlooked in Kaviraj et al. (2012). We also note that a baseline model trained merely on the 352 D-ETGs from Kaviraj et al. (2012) yielded extremely low-purity results, rendering effective detection unfeasible.
@@
Our approach ultimately identifies 520 D - ETGs that were overlooked in Kaviraj et al. (2012), thereby opening up new avenues for understanding merger events and dust properties in nearby galaxies.
5 Discussion
@@ 6 Conclusions
    We fine-tuned the pre-trained conditional diffusion model Stable-Diffusion-v1-5 using high-quality morphology label - galaxy image pairs from the Galaxy Zoo 2 dataset. The fine-tuned model demonstrates excellent performance in synthesizing high-fidelity images that closely follow morphological instructions.
@@
    We further demonstrated that our diffusion model functions as a “world simulator” surrogate, enabling generative extrapolation from well-observed galaxies to rare or even unobserved domains. Using early-type galaxies with prominent dust lanes as a test case (accounting for approximately 0.1% of the Galaxy Zoo 2 dataset and excluded from the training data beforehand), we showed that our generative model can successfully synthesize these galaxies and enhance downstream machine learning models. Our model led to 520 additional dusty early-type galaxies from the same dataset, doubling the sample size compared to previous studies, which only reported 352 samples. This newly created catalog holds scientific value as a proxy for studying merger remnants and dust properties.
@@ 7 Data Availability
The model pipeline in this study is available at https://github.com/chenruiRae/GalaxySD and the trained diffusion model weights are available at https://huggingface.co/CosmosDream/GalaxySD. For any additional information or specific requests regarding the data or code, please contact the corresponding author.
8 Author Contributions
CM conducted experimental runs, created scientific visualizations, and contributed to draft revision. ZS oversaw program design, developed code for the embedding model, and led manuscript writing and revision efforts. TJ gave the initial implementation of the diffusion model fine-tuning, participated in scientific discussions, and provided critical suggestions for manuscript refinement. ZC, YST, and SH offered expert scientific and technical guidance, along with constructive feedback on the manuscript. YST additionally contributed to the selection of the D-ETG research topic and the D-ETG catalog. MYL developed the visual inspection program utilized in the analysis.
9 Acknowledgments
ZS and SH acknowledge support from the National Natural Science Foundation of China (Grants No. 12273015 & 12433003) and the China Crewed Space Program through its Space Application System. ML and ZC acknowledge support from the National Key R&D Program of China (grant no. 2023YFA1605600) and Tsinghua University Initiative Scientific Research Program (No. 20223080023). The authors thank Hongming Tang, Yanhan Guo and Huiling Liu for their valuable discussions and constructive comments. We also appreciate the helpful technical feedback provided by Morph AI Inc. The authors acknowledge the Tsinghua Astrophysics High-Performance Computing platform at Tsinghua University for providing computational and data storage resources that have contributed to the research results reported within this paper.
@@ Appendix A SimCLR as a embedding tool for galaxy image
As our evaluation metric described in Section 4 involves an encoder $f_{\phi}:\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}\rightarrow\mathbb{R}^{\mathrm{L}}$ to map high-dimensional galaxy images into low-dimensional latent representations, we here detail the representation learning technique used in this work – SimCLR here.
SimCLR Chen et al. (2020) is a widely-used self-supervised learning technique in computer vision. The core idea behind, named as constractive learning, is to construct a series positive and negative pair through data augmentation, and then train a neural network to pull together representations of two differently augmented views from the same galaxy image, while simultaneously pushing apart representations from distinct images. This contrastive mechanism drives the encoder to learn invariant features that robustly capture the morphological details of galaxies, thereby creating a highly discriminative embedding space. As a result, the learned representations not only improve the performance in downstream tasks such as image synthesis and classification but also enhance the overall ability to discern subtle astrophysical features. Compared to other representation learning techniques such as masked image modeling (e.g., Xie et al., 2021) and variational autoencoders (e.g., Kingma & Welling, 2013), constractive learning balances both model performance and the cost to training the model and already has wide applications in astrophysics (e.g., Stein et al., 2022; Desmons et al., 2024; Mohale & Lochner, 2024).
Formally, let $\mathbf{x}_{i}$ be an image of a galaxy, and let $\mathbf{x}_{i}^{a}$ and $\mathbf{x}_{i}^{b}$ be two augmented views of $\mathbf{x}_{i}$. These views form a positive pair $(\mathbf{x}_{i}^{a},\mathbf{x}_{i}^{b})$. For a batch of $N$ images, we generate $2N$ augmented views, resulting in $N$ positive pairs and $2N(2N-2)$ negative pairs. The goal is to learn an encoder $f_{\phi}$ that maps images to a latent space where positive pairs are close and negative pairs are far apart.
@@
To implement the SimCLR framework, we utilize a ResNet18 architecture as our encoder $f_{\phi}$. ResNet18, a widely-used convolutional neural network, is chosen for its balance between depth and computational efficiency, making it suitable for extracting rich features from galaxy images. The network is initialized with weights pre-trained on ImageNet to leverage transfer learning benefits. We train the model on the Galaxy Zoo dataset, which includes 300,000 annotated galaxy images from the Sloan Digital Sky Survey (SDSS). Each image undergoes a series of augmentations, such as random cropping, flipping, and color jittering, to generate positive pairs. The model is trained for 200 epochs with a batch size of 256 and a learning rate of 0.001, using the Adam optimizer. The temperature parameter $\tau$ is set to 0.5.
@@ Appendix B Heriarichcal Morphological Labelling System in GZ2
The GZ2 labelling system uses a hierarchical framework starting with broad early/late-type galaxy distinctions, followed by detailed features like bars, dust lanes, and merger signatures. This structured approach enables systematic annotation of both global and fine-grained morphological attributes, as shown in Figure B3.
@@
Appendix C Cross-attention Mechanism in Conditional Diffusion Model
In the diffusion model, conditioning on textual prompts is achieved through a cross-attention mechanism integrated into the U-Net backbone of the denoising process. Instead of conditioning the model via concatenation or direct feature injection, it employs text-image cross-attention mechanism at multiple layers, enabling fine-grained and semantically aligned generation guided by language.
Let $\mathbf{x}_{t}\in\mathbb{R}^{C\times H\times W}$ denote the noisy latent representation at timestep $t$, and let $\boldsymbol{\tau}=[\tau_{1},\ldots,\tau_{n}]$ represent the embedded text tokens, where each $\tau_{i}\in\mathbb{R}^{d}$ is obtained from a frozen text encoder. The cross-attention module computes output features $\mathbf{y}\in\mathbb{R}^{d}$ using the following attention mechanism:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}$ (C1)
In the cross-attention block, the queries $\mathbf{Q}\in\mathbb{R}^{hw\times d}$ are derived from the spatial latent features, while the keys $\mathbf{K}\in\mathbb{R}^{n\times d}$ and values $\mathbf{V}\in\mathbb{R}^{n\times d}$ come from the encoded text tokens $\boldsymbol{\tau}$. This enables spatially-aware modulation of the visual features based on semantic relevance to the text.
The cross-attention is inserted into the U-Net’s ResNet blocks as follows:
$\mathbf{x}_{\text{out}}=\mathbf{x}_{\text{in}}+\text{Attention}(\phi(\mathbf{x}_{\text{in}}),\boldsymbol{\tau},\boldsymbol{\tau})$ (C2)
where $\phi$ denotes a linear projection aligning the latent features to the attention embedding dimension. Multi-head attention Vaswani et al. (2017) is used to capture diverse semantic relationships between image regions and language concepts.
During inference, to enhance targeted morphological features (e.g., dust lanes in this study), we manually assign weights to the corresponding text embeddings to emphasize the desired features, as illustrated in Table 1 and Figure 7. Such enhancement is critical when pushing the boundaries of diffusion models for generative extrapolation tasks.
Appendix D Additional Sanity Check on our D-ETG Samples

### v2

@@ IV.3 Enhancing Early-type Dust Lane Galaxy Detection via Extrapolation on Rare Morphologies
Figure 6: Generative extrapolation of early-type galaxies with prominent dust lanes (D-ETGs). While early-type galaxies are commonly quiescent, our generative model synthesizes D-ETGs by merging the morphology feature in early-type galaxy and star-forming feature (dust lane) in late-type galaxies. Both early-type galaxies and dust lanes are common in the universe, but early-type galaxies with dust lanes are rare since they are origin from recent merging events. Our generation model learns these two concepts from extensive datasets and extrapolates them to synthesize rare D-ETGs, which are excluded in training set. These synthesized D-ETG images are then used as training data for classifiers, facilitating the identification of 520 D-ETGs within a sample of 239,695 GZ2 images as shown in Figure 7.
@@
Figure 7: Examples of D-ETGs synthesized by our diffusion model (upper panel) and newly identified instances using machine learning models augmented with the synthesized data (lower panel). Our diffusion model generates high-fidelity D-ETG images by prompts like “sdss, {completely round:1.2} galaxy, {dust lane:1.3}” or “sdss, in-between round galaxy, {dust lane:1.2}”, where the weights can vary reasonably within the range of 1.1 to 1.4. We manually assign weights to different morphology features through cross-attention mechanism to emphasize generation of desired features as in Section II.
@@
We initiate our investigation with an experimental test. Using the D-ETG catalog from Kaviraj et al. (2012), we aim to verify the effectiveness of our pipeline. In the experimental setup, we randomly select 100 galaxies from the D-ETG catalog in Kaviraj et al. (2012) and 10,000 galaxies from the GZ2 dataset to construct the training set, composing an experimental dataset with 1/100 postive / negative sample ratios, mimicking the data distribution in Kaviraj et al. (2012). Similarly to the approach in Section IV.2, we use pure supervised learning without any additional adjustments as our baseline.
Again, we introduce Focal Loss as an alternative baseline for dealing with imbalance classes as in Section IV.2. This method reweights the contributions of positive and negative samples in the loss function. We note that we defer detailed comparisons with more advanced few - shot algorithms and unsupervised classification algorithms to future studies, as their inclusion is beyond the scope of this work. More importantly, our methodologies are not in competition with these algorithms but can be used together to enhance these algorithms through data augmentation, a topic that we will discuss further in Section V.
@@
Figure 8: Kernel density estimates (KDE) of redshift distributions for D-ETGs identified by our method (red) and Kaviraj et al. (2012) (blue). Redshift values are obtained from the GSWLC Salim et al. (2016). Our sample encompasses relatively higher redshifts, a regime that is more challenging for human volunteers to identify, thereby demonstrating the efficacy of our synthesized image-enhanced machine learning model in detecting rare galaxy populations.
@@
Here, we apply our method to 239,695 Galaxy Zoo 2 (GZ2) images to identify previously undetected early-type dust lane galaxies (D-ETGs). We first construct training data: 10,000 synthetic D-ETGs as positive samples and 10,000 randomly selected normal galaxies as negative samples, to train a binary classifier for D-ETGs. Given the low occurrence of D-ETGs, we assume minimal contamination in the negative sample set. We follow the training procedure described in Section IV.2. After the initial training, the model was applied to the entire GZ2 dataset, identifying 8,184 positive candidates. These candidates were then visually inspected, and the classifier was retrained based on the validated results, yielding 2,488 refined positive samples. This iterative process was repeated once more, resulting in an additional 1,062 positive samples. All candidate samples in the three iteration training stages were visually validated by three authors (CM, ZS, and TJ). Final selection required agreement from at least two reviewers, resulting in a catalog of 520 high-confidence D-ETGs overlooked in Kaviraj et al. (2012). We also note that a baseline model trained merely on the 352 D-ETGs from Kaviraj et al. (2012) yielded extremely low-purity results, rendering effective detection unfeasible.
@@
Our approach ultimately identifies 520 D-ETGs that were overlooked in Kaviraj et al. (2012), thereby opening up new avenues for understanding merger events and dust properties in nearby galaxies. We provide this new sample as an open-access catalog to facilitate further studies of early-type dust lane galaxies, as described in Section VII. The catalog contains, for each of the 520 D-ETGs, the SDSS DR7 object identifier, the corresponding filename as an asset identifier (asset_id), RA, DEC, and subsample identification.
V Discussion
@@ VI Conclusions
    We fine-tuned the pretrained conditional diffusion model Stable-Diffusion-v1-5 using high-quality morphology label - galaxy image pairs from the Galaxy Zoo 2 dataset. The fine-tuned model demonstrates excellent performance in synthesizing high-fidelity images that closely follow morphological instructions.
@@
    We further demonstrated that our diffusion model functions as a “world simulator” surrogate, enabling generative extrapolation from well-observed galaxies to rare or even unobserved domains. Using early-type galaxies with prominent dust lanes as a test case (accounting for approximately 0.1% of the Galaxy Zoo 2 dataset and excluded from the training data beforehand), we showed that our generative model can successfully synthesize these galaxies and enhance downstream machine learning models. Our method led to 520 additional dusty early-type galaxies from the same dataset, doubling the sample size compared to previous studies, which only reported 352 samples. This newly created catalog holds scientific value as a proxy for studying merger remnants and dust properties.
@@ VII Data Availability
The model pipeline in this study is available at the GalaxySD [^6] https://github.com/chenruiRae/GalaxySD with DOI: 10.5281/zenodo.17112014 Github repository. And the trained diffusion model weights [^7] https://huggingface.co/CosmosDream/GalaxySD with DOI: 10.57967/hf/6479 are available as well. For any additional information or specific requests regarding the data or code, please feel free to contact us.
VIII Author Contributions
CM conducted experimental runs, created scientific visualizations, and contributed to draft revision and writing. ZS oversaw program design, developed code for the embedding model, and led manuscript writing and revision efforts. TJ gave the initial implementation of the diffusion model fine-tuning, participated in scientific discussions, and provided critical suggestions for manuscript refinement. ZC, YST, and SH offered expert scientific and technical guidance, along with constructive feedback on the manuscript. YST additionally contributed to the selection of the D-ETG research topic and the D-ETG catalog. MYL developed the visual inspection program utilized in the analysis.
IX Acknowledgments
ZS and SH acknowledge support from the National Natural Science Foundation of China (Grants No. 12273015 & 12433003) and the China Crewed Space Program through its Space Application System. ML and ZC acknowledge support from the National Key R&D Program of China (grant no. 2023YFA1605600) and Tsinghua University Initiative Scientific Research Program (No. 20223080023). CM acknowledge support by the Major Key Project of Peng Cheng Laboratory. The authors thank Hongming Tang, Yanhan Guo and Huiling Liu for their valuable discussions and constructive comments. We also appreciate the helpful technical feedback provided by Morph AI Inc. The authors acknowledge the Tsinghua Astrophysics High-Performance Computing platform at Tsinghua University for providing computational and data storage resources that have contributed to the research results reported within this paper.
@@ Appendix A SimCLR as a embedding tool for galaxy image
As our evaluation metric described in Section IV involves an encoder $f_{\phi}:\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}\rightarrow\mathbb{R}^{\mathrm{L}}$ to map high-dimensional galaxy images into low-dimensional latent representations, we here detail the representation learning technique used in this work – SimCLR here.
SimCLR Chen et al. (2020) is a widely-used self-supervised learning technique in computer vision. The core idea behind, named as contrastive learning, is to construct a series positive and negative pair through data augmentation, and then train a neural network to pull together representations of two differently augmented views from the same galaxy image, while simultaneously pushing apart representations from distinct images. This contrastive mechanism drives the encoder to learn invariant features that robustly capture the morphological details of galaxies, thereby creating a highly discriminative embedding space. As a result, the learned representations not only improve the performance in downstream tasks such as image synthesis and classification but also enhance the overall ability to discern subtle astrophysical features. Compared to other representation learning techniques such as masked image modeling (e.g., Xie et al., 2021) and variational autoencoders (e.g., Kingma & Welling, 2013), contrastive learning balances both model performance and the cost to training the model and already has wide applications in astrophysics (e.g., Stein et al., 2022; Desmons et al., 2024; Mohale & Lochner, 2024).
Formally, let $\mathbf{x}_{i}$ be an image of a galaxy, and let $\mathbf{x}_{i}^{a}$ and $\mathbf{x}_{i}^{b}$ be two augmented views of $\mathbf{x}_{i}$. These views form a positive pair $(\mathbf{x}_{i}^{a},\mathbf{x}_{i}^{b})$. For a batch of $N$ images, we generate $2N$ augmented views, resulting in $N$ positive pairs and 2N(2N2)2N(2N-2) negative pairs. The goal is to learn an encoder $f_{\phi}$ that maps images to a latent space where positive pairs are close and negative pairs are far apart.
@@
To implement the SimCLR framework, we utilize a ResNet18 architecture as our encoder $f_{\phi}$. ResNet18, a widely-used convolutional neural network, is chosen for its balance between depth and computational efficiency, making it suitable for extracting rich features from galaxy images. The network is initialized with weights pretrained on ImageNet to leverage transfer learning benefits. We train the model on the Galaxy Zoo dataset, which includes 300,000 annotated galaxy images from the Sloan Digital Sky Survey (SDSS). Each image undergoes a series of augmentations, such as random cropping, flipping, and color jittering, to generate positive pairs. The model is trained for 200 epochs with a batch size of 256 and a learning rate of 0.001, using the Adam optimizer. The temperature parameter $\tau$ is set to 0.5.
@@ Appendix B Heriarichcal Morphological Labelling System in GZ2
The GZ2 labeling system uses a hierarchical framework starting with broad early/late-type galaxy distinctions, followed by detailed features like bars, dust lanes, and merger signatures. This structured approach enables systematic annotation of both global and fine-grained morphological attributes, as shown in Figure B3.
@@
Appendix C Weighted Prompt of dust lane feature
As we have described in Section II.2, the weights are incorporated into the conditional text embeddings to guide the generation process. This strategy enables us to systematically enhance a target feature (e.g., dust lane) while simultaneously monitoring its impact on other morphological attributes. To quantitatively evaluate and determine the appropriate weights for emphasizing specific features during generation, we conducted controlled experiments in which the weight assigned to the “dust lane” feature was gradually varied. By analyzing how the generated images respond to different weight values, we first show how the weight quantitatively influences the desired feature, as well as how increasing or decreasing this weight affects other correlated features. This analysis provides a principled basis for how to assign reasonable weight values in prompts when emphasizing particular morphological characteristics.
Figure C4: Effect of varying the weight of the “dust lane” feature in the text prompt on generated galaxy image realism scores. The (a) and (b) are two different prompts. In each sub-figure, the top panel shows the tendency of realism scores with respect to the “dust lane” weight (from 1.0 to 1.9) in the prompt word, peaking near 1.2 and then gradually declining as the weight increases from 1.5 to 1.9. The bottom two rows of images are example generated galaxies corresponding to each weight value annotated below. Increasing the weight initially strengthens the presence of dust lanes, but excessively high weights ($\geq$ 1.6) introduce unrealistic textures and artifacts, reducing the perceived realism.
Figure C5: Effect of varying the weight of the “dust lane” and “completely round” feature in the text prompt “{completely round:X} galaxy, {dust lane:Y}” on generated galaxy image realism scores, where X and Y are weights to the corresponding feature. The integration of these two features contributes to the final D-ETG generation. The highest realism scores of generated D-ETG images occur in the region where X $\in[1.1,\,1.4]$ and Y $\in[1.05,\,1.3]$. In the right panel, the boundary is formed where the weights of “completely round” and “dust lane” are equal, as shown by the light yellow boxes and dashed lines. An increase of about +0.1+0.1 weight in “dust lane” feature beyond this boundary contributes to a relatively ideal D-ETGs, as shown by the dark yellow boxes. The farther the weight is from this boundary, the weaker the influence of the other feature in the generated images.
Appendix D Focal Loss Implementation
The focal loss is implemented by first computing the standard cross-entropy loss for each sample without reduction. Let $p_{t}$ denote the predicted probability of the ground-truth class, obtained as
$p_{t}=\exp(-\mathrm{CE}),$
where $\mathrm{CE}$ is the cross-entropy loss.
The final focal loss for each sample is then calculated as
$\mathrm{FL}=\alpha(1-p_{t})^{\gamma}\cdot\mathrm{CE},$
where $\alpha$ is a weighting factor to address class imbalance, and $\gamma$ is the focusing parameter that reduces the relative loss contribution from well-classified examples, thus emphasizing hard-classified examples. Following the parameter settings reported in previous work by Liu et al. (2025b), we adopted the same parameters with $\alpha=0.25$ and $\gamma=4$ to balance minority and majority classes while emphasizing hard-to-classify samples. After parameter adjustments on our dataset, this configuration consistently yielded stable convergence and superior validation performance compared with alternative parameter combinations, confirming its suitability for our task.
The per-sample focal losses are aggregated by taking the mean over the batch, which stabilizes training and ensures that the loss magnitude is independent of the batch size. In our implementation, we set the reduction mode to mean to achieve this effect.
Appendix E Additional Sanity Check on our D-ETG Samples